OpenStudy (anonymous):

. A ball is thrown directly downward with an initial speed of 8.00 m/s, from a height of 30.0 m. After what time interval does it strike the ground?

OpenStudy (anonymous):

it is free fall. so simple accellerated motion. try using the common equation

OpenStudy (anonymous):

So T= sqrt(2(30)/8) ?

OpenStudy (anonymous):

u=u0+1/2at^2 where a=g=10m/s^2

OpenStudy (anonymous):

Ok. Explain further if you can